OpenStudy (anonymous):

(5y+4)(5y-4) ?? solve multiply foil method

OpenStudy (stefrheart):

What is the part that you need help on?

OpenStudy (anonymous):

Foil method.

OpenStudy (fibonaccichick666):

What do the letters of FOIL stand for?

OpenStudy (anonymous):

First Outter Inner Last

OpenStudy (anonymous):

F-first o-outer i-inner l-last

OpenStudy (fibonaccichick666):

good, so now, do you remember the distributive property?

OpenStudy (anonymous):

multiply the first numbers. example 5y*5y=25y(square)

OpenStudy (fibonaccichick666):

foil, is essentially distributing a polynomial to another polynomial

OpenStudy (anonymous):

i need the answser. i know how to do it its just when i did it by hand and cheked on online calculator it came out diffrent

OpenStudy (shinalcantara):

(5y+4)(5y-4) Using FOIL method: Multiply the FIRST terms of the two quantities 5y(5y) = ? ------------ Multiply the OUTER terms of the two quantities 5y(-4) = ? ------------- Multiply the INNER terms of the two quantities 4(5y) = ? ------------- Multiply the LAST terms of the two quantities 4(-4) = ?

OpenStudy (fibonaccichick666):

can you show me your work please @jackelyn_1234 then I can see where you went wrong

OpenStudy (anonymous):

(5y+4)(5y-4) 25y^2-20y+20y+0 25y^2 thats what i got

OpenStudy (anonymous):

25y(square)-40y-16... @shinalcantara is that correct?

OpenStudy (fibonaccichick666):

that way you get it right next time :)

OpenStudy (shinalcantara):

recheck the middle term @megz08

OpenStudy (fibonaccichick666):

no megz, but you may want to check this out www.openstudy.com/code-of-conduct

OpenStudy (fibonaccichick666):

@jackelyn_1234 , how did you get a zero for the last term?

OpenStudy (anonymous):

25y(square)-16?

OpenStudy (anonymous):

4*-4= o no?

OpenStudy (shinalcantara):

Can you recheck the last term @jackelyn_1234 ?

OpenStudy (fibonaccichick666):

nah, what is 4*4 normally?

OpenStudy (anonymous):

16

OpenStudy (fibonaccichick666):

then what happens when you multiply a positive and a negative number?

OpenStudy (fibonaccichick666):

good,

OpenStudy (anonymous):

25y(square)-20y-20y-16= 25y(square)-16 how about this? @jackelyn_1234 @shinalcantara

OpenStudy (fibonaccichick666):

you have the correct answer, but your work is wrong meg, check your signs

OpenStudy (anonymous):

i understand

OpenStudy (anonymous):

i multiplied wrong

OpenStudy (fibonaccichick666):

you just added @jackelyn_1234, but also, you can simplify your answer

OpenStudy (anonymous):

oh. i got it.. 25y(square)-20+20-16

OpenStudy (fibonaccichick666):

you have 25y^2+20y-20y-16, so my question is what is 1+(-1)=? How about 6+(-6)=? Or even (-100)+100=? so how about y+(-y)=?

OpenStudy (fibonaccichick666):

@jackelyn_1234

OpenStudy (anonymous):

sorry about the sign

OpenStudy (fibonaccichick666):

it's ok happens to all of us. But considering I lose 25% of my grade on mistakes like that, I like to make sure others don't

OpenStudy (anonymous):

yeah, so true!! @FibonacciChick666

OpenStudy (anonymous):

help

OpenStudy (anonymous):

Circle the first term in the trinomial. multiply it by each term in the binomial, and place each result in the spaces provided. Repeat the proccess for the second and third term of the trinomial until all of the spaces are filled in. (3x^3-2x^2+7)(x^3+2x) ___+____+____+_____+____+____

OpenStudy (fibonaccichick666):

make a new question so it's cleaner please

OpenStudy (anonymous):

thats what the problem says i cant put in in any words

OpenStudy (fibonaccichick666):

a new post on this site I mean, so that we don't have to scroll down through all the other posts

OpenStudy (anonymous):

ohh ok